# Strong order-one convergence of the Euler method for random ordinary differential equations driven by semi-martingale noises

### v7

@@ 8.3.8. A random Fisher-KPP nonlinear PDE driven by boundary noise
The random process $\{Y_{t}\}_{t}$, which drives the flux on the left boundary point, is taken to be a colored noise modulated by an exponentially decaying Hawkes process, representing random wave trains of incoming populations. More precisely, $Y_{t}=H_{t}O_{t}$, where $\{H_{t}\}_{t}$ is a Hawkes process with initial rate $\lambda_{0}=5.0$, base rate $a=1.4$, exponential decay rate $\delta=8.0$, and jump law following an exponential distribution with scale $\theta=1/0.4;$ and $\{O_{t}\}_{t}$ is an Ornstein-Uhlenbeck process with $O_{0}=0.0,$ time-scale $\tau=0.005$, drift $\nu=1/\tau$ and diffusion $\sigma=\zeta/\tau,$ where $\zeta=\sqrt{\tau}.$ Since both $\{H_{t}\}_{t}$ and $\{O_{t}\}_{t}$ are semi-martingales and the class of semi-martingales is an algebra [32, Corollary II.3], the modulated process $\{Y_{t}\}_{t}$ is also a semi-martingale.
@@
The equation involves a nonlinear term which is not globally Lipschitz continuous, but, similiarly to the population dynamics model considered in Section 8.3.4, the region $0\leq u(t,x)\leq u_{m}$ is invariant, so that the nonlinear term can be modified outside this region in order to satisfy the required uniform global estimates without affecting the dynamics within this region. The initial condition is chosen to be within this region almost surely. The procedure is the same as that done in Section 8.3.4, so the details are omited.
For the time-mesh parameters, we set $N_{\textrm{tgt}}=2^{18}$ and $N_{i}=2^{5},2^{7},2^{9}.$ The spatial discretization is done with finite differences, with the number of spatial points depending on the time mesh, for stability and convergence reasons. Indeed, the Von Neumann stability analysis requires that $2\mu\Delta t/\Delta_{x}^{2}\leq 1$ With that in mind, for each $N_{i}=2^{5},2^{7},2^{9}$, we take the $K_{i}+1$ spatial points $0=x_{0}<\ldots x_{K_{i}},$ with $K_{i}=2^{3},$ $2^{4},$ and $2^{5},$ respectively, while for the target solution, we use $K_{\textrm{tgt}}=2^{9}.$
@@
Table 9. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for the Fisher-KPP equation for each mesh resolution $N$, with initial condition $X_{0}=0$ and Hawkes-modulated Ornstein-Uhlenbeck colored noise, on the time interval $I=[0.0,2.0]$, based on $M=40$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.116$, with the 95% confidence interval $[0.9995,1.2318]$.
Figure 23. Order of convergence of the strong error of the Euler method for the Fisher-KPP model Equation 8.28-Equation 8.29, based on Table 9.

### v8

@@ A.2.7. A random Fisher-KPP nonlinear PDE driven by boundary noise
$\frac{\partial u}{\displaystyle\partial t}=\mu\frac{\partial^{2}u}{\partial x^{2}}+\lambda u\left(1-\frac{u}{u_{m}}\right),\quad(t,x)\in(0,\infty)\times(0,1),$ (A.26)
@@
$\frac{\partial u}{\partial x}(t,0)=-Y_{t},\quad\frac{\partial u}{\partial x}(t,1)=0,$ (A.27)
@@
The random process $\{Y_{t}\}_{t}$, which drives the flux on the left boundary point, is taken to be a colored noise modulated by an exponentially decaying Hawkes process, representing random wave trains of incoming populations. More precisely, $Y_{t}=H_{t}O_{t}$, where $\{H_{t}\}_{t}$ is a Hawkes process with initial rate $\lambda_{0}=5.0$, base rate $a=1.4$, exponential decay rate $\delta=8.0$, and jump law following an exponential distribution with scale $\theta=1/0.4;$ and $\{O_{t}\}_{t}$ is an Ornstein-Uhlenbeck process with $O_{0}=0.0,$ time-scale $\tau=0.005$, drift $\nu=1/\tau$ and diffusion $\sigma=\zeta/\tau,$ where $\zeta=\sqrt{\tau}.$ Since both $\{H_{t}\}_{t}$ and $\{O_{t}\}_{t}$ are semi-martingales and the class of semi-martingales is an algebra [41, Corollary II.3], the modulated process $\{Y_{t}\}_{t}$ is also a semi-martingale.
@@
The equation involves a nonlinear term which is not globally Lipschitz continuous, but, similiarly to the population dynamics model considered in Section A.2.3, the region $0\leq u(t,x)\leq u_{m}$ is invariant, so that the nonlinear term can be modified outside this region in order to satisfy the required uniform global estimates without affecting the dynamics within this region. The initial condition is chosen to be within this region almost surely. The procedure is the same as that done in Section A.2.3, so the details are omited.
For the time-mesh parameters, we set $N_{\textrm{tgt}}=2^{18}$ and $N_{i}=2^{5},2^{7},2^{9}.$ The spatial discretization is done with finite differences, with the number of spatial points depending on the time mesh, for stability and convergence reasons. Indeed, the Von Neumann stability analysis requires that $2\mu\Delta t/\Delta_{x}^{2}\leq 1.$ With that in mind, for each $N_{i}=2^{5},2^{7},2^{9}$, we take the $K_{i}+1$ spatial points $0=x_{0}<\ldots x_{K_{i}},$ with $K_{i}=2^{3},$ $2^{4},$ and $2^{5},$ respectively, while for the target solution, we use $K_{\textrm{tgt}}=2^{9}.$
@@
Table 9. Mesh points (N), time steps (dt), strong error (error), and standard error (std err) of the Euler method for the Fisher-KPP equation for each mesh resolution NN, with initial condition $X_{0}=0$ and Hawkes-modulated Ornstein-Uhlenbeck colored noise, on the time interval $I=[0.0,2.0]$, based on $M=40$ sample paths for each fixed time step, with the target solution calculated with $262144$ points. The order of strong convergence is estimated to be $p=1.116$, with the 95% confidence interval $[0.9995,1.2318]$.
Figure 23. Order of convergence of the strong error of the Euler method for the Fisher-KPP model Equation A.26-Equation A.27, based on Table 9.